Status of proto-proto:
- 4 fiber-bundles ordered; 3 of them have 2m inner length and 4m outer length; one bundle contains 76 fibers and is thus sufficient for proto-proto; available in 2 months;
- receivers: crate ordered, receivers ready to be tested;
- steering module will arrive at LNGS by mid February
- first MB ok (but SiPM single dose, low uniformity)
- second MB delayed due to late arrival fo SiPMs from FBK run #4 (now expected in first week of February); it seems still possible to have the 2nd MB ready by end of February; the SiPMs are triple dose, high uniformity
- TPC: Clevios painting received at CERN; tests will start next week (mechanical resistance, electrical conductivity); tests on two sided wires to be performed at Houston will be reported at next meeting; updated design to be finalized only after test completion; once it is available it must be sent to Bologna in order to finalize tools for MB;
- test set-up at CERN: in Francesco's lab; HV feed through to be collected at LNGS; purification loop ok
- DAQ: VME crate ok; 5 V1725 boards in TRIUMF; optical interface being shipped; TRIUMF will provide DAQ server, need to organize shipment; order placed for analysis server and disk storage; TRIUMF also finalizing trigger box; MIDAS can interface to slow control;
- need to cover slow control task
- Power supplies and filter boxes not yet available (Alessandro R.)
Next meeting in one week.
Next to next will be a face-to-face meeting at CERN (date tbd January 24 or 25 )
There are minutes attached to this event.